LECTURE SCHEDULE, MAT 123, Spring 2003

Week of  Sections  Holidays and Exams

1/22  2.1
1/27  2.2, 2.3
2/3  2.4, 2.5, 2.6
2/10  2.7, 3.1
2/17  3.6, 4.1, 4.2
2/24  4.3, 4.4
3/3 4.5 First Exam, Mon Mar 3, 8:30 pm
3/10  5.1, 5.2
3/17  No class this week
3/24  5.3, 5.4
3/31  7.1, 7.2
4/7  7.4  Second Exam, Wed Apr 9 8:30 pm
4/14  C2.1  No class Wed-Fri
4/21  C2.2, C2.3
4/28  C2.6, C2.7
5/5 C2.8 Fri May 9 is last day of classes
5/12  Final, Thur May 15, 11-1:30

Rooms for the midterms and finals will be announced in class.

HOMEWORK, MAT 123, Spring 2003 REVISED

Section: Topic  Problems:
2.1  What is a function?  2,4,8,12,28,30,36,50,54
2.2  Graphs of functions  2,4,10,12,18,26,28,40,60
2.3  Applied functions: variation  2,4,8,14,18,22,24,28
2.4  Average rate of change  4,8,12,16,18,20,24
2.5  Transformation of functions  4,8,12,20,22,30,34
2.6  Extreme values of functions  6,16,18,22,24,28,36
2.7  Modeling with functions  4,8,12,18,20,22,26
3.1  Polynomial functions and graphs  10,12,20,58,60,74
3.6  Rational functions  4,8,16,28,34,50
4.1  Exponential functions  4,8,12,28,32,38,42,54
4.2  Logarithmic functions  6,8,12,26,28,36,42,44,58
4.3  Laws of logarithms  4,6,8,12,18,24,36,40,48
4.4  Exponential and logarithmic equations 1,14,20,32,38,46,56,68
4.5  Modeling with exponentials  2,4,8,14,22,24
5.1  The unit circle  6,8,14,18,20,28,38
5.2  Trigonometric functions  6,14,28,36,42,46
5.3  Trigonometric graphs  6,10,20,28,36,50
5.4  More trigonometric graphs  4,12,22,42
7.1  Trigonometric identities  2,4,14,24,34,38,58
7.2  Addition formula  8,16,24,42
7.4  Inverse Trigonometric functions  6,12,14,30,40,48
C2.1  Tangents and velocity  2,4,6
C2.2  Limit of a function  4,8,10,18
C2.3  Limit laws  2,4,12,24,30,38
C2.6  Rates of change  6,16,18,20
C2.7  Derivatives  2,8,16,28
C2.8  Deriviative function  2,6,8,10,24,36
C2.9  Linear approximations  8,10,14
C2.10  What does f′ say about f?  2,8,12,16,18,26,28

Sections starting with “C” refer to the calculus part of the text.